Change page style: 

OT Library

The Observing Tool libraries contain "pieces" of observations e.g. commonly used configurations, sequences, standard calibrations and ancillary information which may be copy-and-pasted into a new observation. The intent of these libraries is to speed the development of new science programs. As observing procedures and instrumental configurations may change from semester to semester (installation of new GCAL filters, for instance), it is advisable that PIs download the most recent libraries when defining new observations.

Announcement to T-ReCS Users in Semester 2002A

Phil Puxley, Gemini Associate Director

22 February 2002

Progress is being made at the University of Florida towards preparing T-ReCS for pre-ship acceptance tests prior to its delivery to Gemini. Although substantially improved, issues remain with flexure and elevated system noise. In addition, several key measurements such as the internal cryostat background and characterisation of the science-grade detector have yet to be carried out.

T-ReCS Mounted on Gemini South

13 May 2003

After three weeks of integration and testing on Cerro Pachon, T-ReCS was successfully mounted on the Gemini South telescope today. The initial on-telescope test results show the same detector and optical characteristics as measured in the lab. First-light is scheduled for early June.

T-ReCS Completes Imaging Commissioning and begins System Verification Science Observations

16 September 2003

Commissioning of the principal imaging mode of T-ReCS was completed during the most recent run in September and the first science data for several System Verification (SV) programs have been obtained. This milestone was the culmination of several runs in June, July and August, often fighting against poor winter weather (e.g. 15 nights out of the first 20 unusable for mid-IR observing). 

Two Issues With T-ReCS Optical Performance

5 March 2005

Two recent issues regarding T-ReCS optical performance have recently come to our attention.

Filter and Window Changes; Implementation of an Occulting Bar

24 January 2008

Filter Red Leaks Fixed

The problems discussed in the 5 March 2005 Hot News concerning the red leaks in the Si-3 9.7 um and Si-4 10.4 um filters have been fixed. The old BaF2 blocking elements were replaced in November 2005 with CaF2 elements and have since been shown to be reliable in removing any light leaks.